OpenStudy (anonymous):

Factor: 10s^2 - 74s - 48

OpenStudy (anonymous):

Use this variation on the quadratic formula to help you factor painful quadratics: \[ax^2+bx+c = a\left(x + \frac{b+\sqrt{b^2-4ac}}{2a}\right)\left(x+\frac{b-\sqrt{b^2-4ac}}{2a}\right)\]As an exercise, I'd suggest you distribute the terms on the right-hand side to confirm the identity.